# ARM: Refining Multivariate Forecasting with Adaptive Temporal-Contextual Learning

### v1

@@ 4 Experiments and Results
We have conducted extensive experiments on 9 widely-used datasets, including the ETT(Zhou et al., 2021), Traffic, Electricity, Weather, ILI, and Exchange Rate(Lai et al., 2018) datasets (details in §A.1). We also generated a dataset ”Multi,” in which we employed a simple shifting operation to construct significant inter-series relationships (see Figure 1 (c), Figure 5, and §A.2). We considered 6 baselines for comparison. Our selection included Transformer-based models such as PatchTST(Nie et al., 2023), FEDformer(Zhou et al., 2022), Autoformer(Wu et al., 2022), and Informer(Zhou et al., 2021), as well as a linear model DLinear from (Zeng et al., 2022). Additionally, we included a last-value ”Repeat” method for comparison. All baselines adhered to the same experimental setup and have 4 different $L_{P}$. We report the baseline results by referring to their original papers. For models that did not report their performance on specific datasets, we conducted the experiments correspondingly. The performance was evaluated using mean squared error (MSE) and mean absolute error (MAE). The hyperparameter settings and implementation are detailed in §A.4 and §A.5.
In Table 1, we evaluate the performance of ARM across 10 datasets. Note that the ARM here employs a vanilla Transformer encoder-decoder, as illustrated in Figure 3, which is referred to as ”Vanilla”. The results demonstrate that ARM consistently outperforms previous models across 10 benchmarks. Its accurately modelling of multivariate relationships leads to enhanced performance on datasets with strong inter-series relationships and larger series scales, like Electricity, and Traffic. Also, extracting univariate effects allows better performance on datasets with short-term distribution changes like Exchange and ILI. ARM efficiently handles datasets of different scales and series-wise relationship intensities by extracting univariate effects, building suitable local representations, and preventing overfitting from incorrect inter-series patterns, thus adeptly manages the characteristic differences in multivariate input. Building on the vanilla Transformer and only slightly adding the computational costs (see §A.6), ARM offers adaptation across diverse multivariate datasets.
Table 1: Multivariate results with prediction horizon $L_{P}\in\{96,192,336,720\}$ (and $L_{P}\in\{24,36,48,60\}$ for ILI). The best results are highlighted bold and the second best are underlined. Here, ARM, in vanilla settings as depicted in Figure 3, employs the vanilla Transformer as its encoder-decoder. Results for other models are sourced from the original literature or derived from additional experiments conducted for unreported datasets. Other models were tested at $L_{I}$ values of ${96,192,336,720}$, and best results among them are reported. In contrast, We use a consistent $L_{I}$ of 720 for ARM.
@@
Electricity 96 0.125 0.222 0.129 0.222 0.140 0.237 0.193 0.308 0.201 0.317 0.274 0.368 1.588 0.946
192 0.142 0.239 0.147 0.240 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 1.595 0.950
336 0.154 0.251 0.163 0.259 0.169 0.267 0.214 0.329 0.231 0.338 0.300 0.394 1.617 0.961
720 0.179 0.275 0.197 0.290 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 1.647 0.975
ETTm1 96 0.287 0.340 0.293 0.346 0.299 0.343 0.326 0.390 0.510 0.492 0.626 0.560 1.214 0.665
192 0.328 0.364 0.333 0.370 0.335 0.365 0.365 0.415 0.514 0.495 0.725 0.619 1.261 0.690
336 0.364 0.384 0.369 0.392 0.369 0.386 0.392 0.425 0.510 0.492 1.005 0.741 1.283 0.707
720 0.411 0.412 0.416 0.420 0.425 0.421 0.446 0.458 0.527 0.493 1.133 0.845 1.319 0.729
ETTm2 96 0.163 0.254 0.166 0.256 0.167 0.260 0.203 0.287 0.255 0.339 0.365 0.453 0.266 0.328
192 0.218 0.290 0.223 0.296 0.224 0.303 0.269 0.328 0.281 0.340 0.533 0.563 0.340 0.371
336 0.265 0.324 0.274 0.329 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 0.412 0.410
720 0.357 0.382 0.362 0.385 0.397 0.421 0.421 0.415 0.422 0.419 3.379 1.388 0.521 0.465
ETTh1 96 0.366 0.391 0.370 0.400 0.375 0.399 0.376 0.415 0.435 0.446 0.941 0.769 1.295 0.713
192 0.402 0.421 0.413 0.429 0.405 0.416 0.423 0.446 0.456 0.457 1.007 0.786 1.325 0.733
336 0.421 0.431 0.422 0.440 0.439 0.443 0.444 0.462 0.486 0.487 1.038 0.784 1.323 0.744
720 0.437 0.459 0.447 0.468 0.472 0.490 0.469 0.492 0.515 0.517 1.144 0.857 1.339 0.756
ETTh2 96 0.264 0.327 0.274 0.337 0.289 0.353 0.332 0.374 0.332 0.368 1.549 0.952 0.432 0.422
192 0.327 0.374 0.341 0.382 0.383 0.418 0.407 0.446 0.426 0.434 3.792 1.542 0.534 0.473
336 0.356 0.393 0.329 0.384 0.448 0.465 0.400 0.447 0.477 0.479 4.215 1.642 0.591 0.508
720 0.371 0.408 0.379 0.422 0.605 0.551 0.412 0.469 0.453 0.490 3.656 1.619 0.588 0.517
Weather 96 0.144 0.193 0.149 0.198 0.176 0.237 0.217 0.296 0.266 0.336 0.300 0.384 0.259 0.254
192 0.189 0.240 0.194 0.241 0.220 0.282 0.276 0.336 0.307 0.367 0.598 0.544 0.309 0.292
336 0.232 0.280 0.245 0.282 0.265 0.319 0.339 0.380 0.359 0.395 0.578 0.523 0.377 0.338
720 0.296 0.332 0.314 0.334 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 0.465 0.394
Traffic 96 0.356 0.247 0.360 0.239 0.410 0.282 0.587 0.366 0.613 0.388 0.719 0.391 2.723 1.079
192 0.373 0.258 0.379 0.246 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 2.756 1.087
336 0.383 0.274 0.392 0.264 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.420 2.791 1.095
720 0.425 0.294 0.432 0.286 0.466 0.315 0.626 0.382 0.660 0.408 0.864 0.472 2.811 1.097
Exchange 96 0.078 0.197 0.087 0.207 0.081 0.203 0.148 0.278 0.197 0.323 0.847 0.752 0.081 0.196
192 0.150 0.280 0.194 0.316 0.157 0.293 0.271 0.380 0.300 0.369 1.204 0.895 0.167 0.289
336 0.252 0.367 0.351 0.432 0.305 0.414 0.460 0.500 0.509 0.524 1.672 1.036 0.305 0.396
720 0.486 0.535 0.867 0.697 0.643 0.601 1.195 0.841 1.447 0.941 2.478 1.310 0.823 0.681
ILI 24 1.148 0.699 1.319 0.754 2.215 1.081 3.228 1.260 3.483 1.287 5.764 1.677 6.587 1.701
36 1.352 0.783 1.579 0.870 1.963 0.963 2.679 1.080 3.103 1.148 4.755 1.467 7.130 1.884
48 1.497 0.799 1.553 0.815 2.130 1.024 2.622 1.078 2.669 1.085 4.763 1.469 6.575 1.798
60 1.378 0.771 1.470 0.788 2.368 1.096 2.857 1.157 2.770 1.125 5.264 1.564 5.893 1.677
Multi 96 0.032 0.125 0.072 0.202 0.067 0.190 0.117 0.261 0.162 0.313 0.092 0.219 0.068 0.189
192 0.063 0.173 0.167 0.294 0.137 0.269 0.204 0.330 0.356 0.459 0.207 0.338 0.143 0.273
336 0.164 0.286 0.314 0.405 0.238 0.355 0.313 0.402 0.572 0.705 0.284 0.414 0.264 0.369
720 0.450 0.503 0.700 0.588 0.476 0.522 0.580 0.544 0.705 0.621 0.921 0.795 0.617 0.551
@@
The three core modules of ARM can be easily integrated into existing LTSF models without changing their structure. Replacing the encoder-decoder predictor in Figure 3 with previous models to apply ARM can effectively help them handle multivariate characteristic differences (see §A.5.2 for more details on applying ARM to existing models). The first part of Table 2 shows the performance boost ARM offers to both univariate (PatchTST, DLinear) and multivariate (Vanilla, Autoformer, Informer) models. We used 2 datasets for experiments: the large-scale Electricity dataset with pronounced inter-series relationships and regular temporal patterns, and the smaller ETTm1 dataset with less evident inter-series links and irregular patterns. With a fixed $L_{I}=720$, our experiments show ARM’s consistent performance advantage without the need of tuning $L_{I}$ like previous models. For datasets with weaker inter-series dependencies, like ETTm1, models employing univariate predictors remain slightly superior. Conversely, for datasets with strong dependencies like Electricity, multivariate predictors have a slight edge. The balanced structure of the vanilla Transformer, without over-design, performs consistently well comparing to other models after integrating ARM.

### v2

@@ 4 Experiments and Results
We have conducted extensive experiments on 9 widely-used datasets, including the ETT(Zhou et al., 2021), Traffic, Electricity, Weather, ILI, and Exchange Rate(Lai et al., 2018) datasets (details in §A.1). We also generated a dataset ”Multi,” in which we employed a simple shifting operation to construct significant inter-series relationships (see Figure 1 (c), Figure 6, and §A.2). Our baseline comparison included Transformer-based models such as PatchTST(Nie et al., 2022), FEDformer(Zhou et al., 2022), Autoformer(Wu et al., 2021), and Informer(Zhou et al., 2021), and a linear model DLinear from (Zeng et al., 2023). Additionally, we included a last-value ”Repeat” method for comparison. We report the baseline results by referring to their original papers. For models that did not report their performance on specific datasets, we conducted the experiments correspondingly. The performance was evaluated using mean squared error (MSE) and mean absolute error (MAE). The hyperparameter settings and implementation are detailed in §A.5 and §A.6.
In Table 1, we evaluate the performance of ARM across 10 datasets. Note that the ARM here employs a vanilla Transformer encoder-decoder, as illustrated in Figure 2, which is referred to as ”Vanilla”. The results demonstrate that ARM consistently outperforms previous models across 10 benchmarks. Its accurately modelling of multivariate relationships leads to enhanced performance on datasets with strong inter-series relationships and larger series scales, like Electricity, and Traffic. Also, extracting univariate effects allows better performance on datasets with short-term distribution changes like Exchange and ILI. Building on the vanilla Transformer and only slightly adding the computational costs (see §A.7), ARM offers adaptation across diverse multivariate datasets.
Table 1: Multivariate results with prediction horizon $L_{P}\in\{96,192,336,720\}$ (and $L_{P}\in\{24,36,48,60\}$ for ILI). The best results are highlighted bold and the second best are underlined. Results for baseline models are sourced from previous literature or derived from additional experiments conducted for unreported datasets. Baselines were run with $L_{I}$ values of 96,192,336,720{96,192,336,720}, and best results among them are reported. In contrast, We use a consistent $L_{I}$ of 720 for ARM.
@@
Electricity (96) 0.125 0.222 0.129 0.222 0.140 0.237 0.193 0.308 0.201 0.317 0.274 0.368 1.588 0.946
Electricity (192) 0.142 0.239 0.147 0.240 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 1.595 0.950
Electricity (336) 0.154 0.251 0.163 0.259 0.169 0.267 0.214 0.329 0.231 0.338 0.300 0.394 1.617 0.961
Electricity (720) 0.179 0.275 0.197 0.290 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 1.647 0.975
ETTm1 (96) 0.287 0.340 0.293 0.346 0.299 0.343 0.326 0.390 0.510 0.492 0.626 0.560 1.214 0.665
ETTm1 (192) 0.328 0.364 0.333 0.370 0.335 0.365 0.365 0.415 0.514 0.495 0.725 0.619 1.261 0.690
ETTm1 (336) 0.364 0.384 0.369 0.392 0.369 0.386 0.392 0.425 0.510 0.492 1.005 0.741 1.283 0.707
ETTm1 (720) 0.411 0.412 0.416 0.420 0.425 0.421 0.446 0.458 0.527 0.493 1.133 0.845 1.319 0.729
ETTm2 (96) 0.163 0.254 0.166 0.256 0.167 0.260 0.203 0.287 0.255 0.339 0.365 0.453 0.266 0.328
ETTm2 (192) 0.218 0.290 0.223 0.296 0.224 0.303 0.269 0.328 0.281 0.340 0.533 0.563 0.340 0.371
ETTm2 (336) 0.265 0.324 0.274 0.329 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 0.412 0.410
ETTm2 (720) 0.357 0.382 0.362 0.385 0.397 0.421 0.421 0.415 0.422 0.419 3.379 1.388 0.521 0.465
ETTh1 (96) 0.366 0.391 0.370 0.400 0.375 0.399 0.376 0.415 0.435 0.446 0.941 0.769 1.295 0.713
ETTh1 (192) 0.402 0.421 0.413 0.429 0.405 0.416 0.423 0.446 0.456 0.457 1.007 0.786 1.325 0.733
ETTh1 (336) 0.421 0.431 0.422 0.440 0.439 0.443 0.444 0.462 0.486 0.487 1.038 0.784 1.323 0.744
ETTh1 (720) 0.437 0.459 0.447 0.468 0.472 0.490 0.469 0.492 0.515 0.517 1.144 0.857 1.339 0.756
ETTh2 (96) 0.264 0.327 0.274 0.337 0.289 0.353 0.332 0.374 0.332 0.368 1.549 0.952 0.432 0.422
ETTh2 (192) 0.327 0.374 0.341 0.382 0.383 0.418 0.407 0.446 0.426 0.434 3.792 1.542 0.534 0.473
ETTh2 (336) 0.356 0.393 0.329 0.384 0.448 0.465 0.400 0.447 0.477 0.479 4.215 1.642 0.591 0.508
ETTh2 (720) 0.371 0.408 0.379 0.422 0.605 0.551 0.412 0.469 0.453 0.490 3.656 1.619 0.588 0.517
Weather (96) 0.144 0.193 0.149 0.198 0.176 0.237 0.217 0.296 0.266 0.336 0.300 0.384 0.259 0.254
Weather (192) 0.189 0.240 0.194 0.241 0.220 0.282 0.276 0.336 0.307 0.367 0.598 0.544 0.309 0.292
Weather (336) 0.232 0.280 0.245 0.282 0.265 0.319 0.339 0.380 0.359 0.395 0.578 0.523 0.377 0.338
Weather (720) 0.296 0.332 0.314 0.334 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 0.465 0.394
Traffic (96) 0.356 0.247 0.360 0.249 0.410 0.282 0.587 0.366 0.613 0.388 0.719 0.391 2.723 1.079
Traffic (192) 0.373 0.258 0.379 0.256 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 2.756 1.087
Traffic (336) 0.383 0.274 0.392 0.264 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.420 2.791 1.095
Traffic (720) 0.425 0.294 0.432 0.286 0.466 0.315 0.626 0.382 0.660 0.408 0.864 0.472 2.811 1.097
Exchange (96) 0.078 0.197 0.087 0.207 0.081 0.203 0.148 0.278 0.197 0.323 0.847 0.752 0.081 0.196
Exchange (192) 0.150 0.280 0.194 0.316 0.157 0.293 0.271 0.380 0.300 0.369 1.204 0.895 0.167 0.289
Exchange (336) 0.252 0.367 0.351 0.432 0.305 0.414 0.460 0.500 0.509 0.524 1.672 1.036 0.305 0.396
Exchange (720) 0.486 0.535 0.867 0.697 0.643 0.601 1.195 0.841 1.447 0.941 2.478 1.310 0.823 0.681
ILI (24) 1.148 0.699 1.319 0.754 2.215 1.081 3.228 1.260 3.483 1.287 5.764 1.677 6.587 1.701
ILI (36) 1.352 0.783 1.579 0.870 1.963 0.963 2.679 1.080 3.103 1.148 4.755 1.467 7.130 1.884
ILI (48) 1.497 0.799 1.553 0.815 2.130 1.024 2.622 1.078 2.669 1.085 4.763 1.469 6.575 1.798
ILI (60) 1.378 0.771 1.470 0.788 2.368 1.096 2.857 1.157 2.770 1.125 5.264 1.564 5.893 1.677
Multi (96) 0.032 0.125 0.072 0.202 0.067 0.190 0.117 0.261 0.162 0.313 0.092 0.219 0.068 0.189
Multi (192) 0.063 0.173 0.167 0.294 0.137 0.269 0.204 0.330 0.356 0.459 0.207 0.338 0.143 0.273
Multi (336) 0.164 0.286 0.314 0.405 0.238 0.355 0.313 0.402 0.572 0.705 0.284 0.414 0.264 0.369
Multi (720) 0.450 0.503 0.700 0.588 0.476 0.522 0.580 0.544 0.705 0.621 0.921 0.795 0.617 0.551
@@
The three core modules of ARM can be easily integrated into existing LTSF models without changing their structure. Replacing the encoder-decoder predictor in Figure 2 with previous models to apply ARM can effectively help them handle multivariate characteristic differences (see §A.6.2 for more details on applying ARM to existing models). The second part of Table 2 shows the improvement ARM offers to both univariate (PatchTST, DLinear) and multivariate (Vanilla, Autoformer, Informer) models. We used 2 datasets for experiments: the large-scale Electricity dataset with pronounced inter-series relationships and regular temporal patterns, and the smaller ETTm1 dataset with less evident inter-series links and irregular patterns. With a fixed $L_{I}=720$, our experiments show ARM’s consistent performance boost without the need of tuning $L_{I}$ like previous models. For datasets with weaker inter-series dependencies, like ETTm1, models employing univariate predictors remain slightly superior. Conversely, for datasets with strong dependencies like Electricity, multivariate predictors have a slight edge. The balanced structure of the vanilla Transformer, without over-design, performs consistently well comparing to other models after integrating ARM.